what function do you know from calculus is such that its first derivative is itself? Its first derivative is a constant multiple k of itself? Write each answer in the form of a first-order differential equation with a solution.


1
Expert's answer
2021-08-30T16:34:33-0400

1.


"\dfrac{dy}{dx}=y"

"\dfrac{dy}{y}=dx"

Integrate


"\int\dfrac{dy}{y}=\int dx"

"\ln|y|=x+\ln C"

"y=Ce^x"


2.


"\dfrac{dy}{dx}=ky"

"\dfrac{dy}{y}=kdx"

Integrate


"\int\dfrac{dy}{y}=\int kdx"

"\ln|y|=kx+\ln C"

"y=Ce^{kx}"